Solve for what is asked in the following problem, show your complete solution



Student A investigated the rate of reaction between calcium carbonate (marble chips) and hydrochloric acid. The student collected 10.0 cm3 of gas produced after 15 seconds.Student B investigated the rate of reaction between magnesium and hydrochloric acid. The student collected 19.5 cm3 of gas produced after 10 seconds.Which student’s investigation had a faster rate of reaction?

1
Expert's answer
2022-04-21T02:04:05-0400

First rate: 10.0 / 15 = 0.667 cm3/s

Second rate: 19.5 / 10 = 1.95 cm3/s

Second / first = 1.95 / 0.667 = 2.9 times greater